OpenStudy (anonymous):

Vector Problem: In triangle ABC, let vector u = AB and vector v = BC. Let M be the midpoint of AB and N be the midpoint of BC. Write each of the following in terms of u and v. a. BA b. CA c. MN

OpenStudy (anonymous):



OpenStudy (phi):

use a - to change the direction of a vector. -AB = BA you can see that AB + BC gets you from A to C, that is, AB+BC is AC you can get from M to N by MB and then BN MB is 1/2 AB and MN is 1/2 BC

OpenStudy (anonymous):

In terms of vectors u and v (u = AB, v = BC): a. BA = -u b. CA = -v - u c. MN = (u/2) + (v/2) Do these answers seem correct?

OpenStudy (phi):

yes, they look good. of course, you might see (obvious) variations CA = -(u+v) MN= (u+v)/2

OpenStudy (anonymous):

Right, that would make it simpler. Thanks for your help